OpenStudy (anonymous):

Use the quadratic formula to solve the equation. x^2-x=-2 Please help, I dont' know why this one is so confusing to me, but I can't get it.

OpenStudy (callisto):

x^2-x=-2 x^2-x+2 =0 ax^2 +bx +c =0 \[x=(-b \pm \sqrt{b^2-4ac)}/(2a)\]

OpenStudy (callisto):

put the corresponding value to the formula

OpenStudy (anonymous):

x^2-x+2=0 Use quadratic formula: ax^2-bx+2=0 a=1, b=-1, c=2

OpenStudy (anonymous):

so the x's are equal to 1 right?

OpenStudy (callisto):

since Δ<0, you will not get real solutions..

OpenStudy (callisto):

i is the imaginary number\[i=\sqrt{-1}\]

OpenStudy (anonymous):

so was my answer close? I thought I was completely wrong and deleted it.

OpenStudy (callisto):

Hmm.. close but not correct... don't be afraid to make mistakes!

OpenStudy (anonymous):

Sorry this is really confusing to me.

OpenStudy (callisto):

So, can you try again to get the answer?

OpenStudy (anonymous):

I can try, but I don't understand still how it works. With no real numbers, it totally confuses me.

OpenStudy (callisto):

show you steps please~

OpenStudy (anonymous):

trying to go over every thing you have said so far to see if I can comprehend it.

OpenStudy (callisto):

i deleted my answer.. i hope you can try...perhaps you can start with substituting the numbers into the formula and ask whatever you don't understand so we can help

OpenStudy (anonymous):

I don't understand the imaginary number. Where does that come in?

OpenStudy (anonymous):

what is the formula I need to use? Is it the ax^2+bx+2?

OpenStudy (callisto):

ax^2 +bx +c =0 is the general form of a quadratic equation. Before you apply the quadratic formula, you have to make the equation into general form first

OpenStudy (callisto):

then use the quadratic formula \[x=(-b \pm \sqrt{b^2-4ac)}/(2a)\]

OpenStudy (anonymous):

Ok now my nerves are gone. I have gone beyond confusion now. LOL

OpenStudy (callisto):

that means you understand it now?

OpenStudy (anonymous):

so when you say to put it into general form...you mean its not already in general form the way that it is?

OpenStudy (anonymous):

No that means that confusion is an understatement. Now I am just plain lost!

OpenStudy (callisto):

x^2-x=-2 it is not in a general form

OpenStudy (callisto):

but when you add 2 on both sides, it is in the general form

OpenStudy (anonymous):

ok

OpenStudy (callisto):

so you 'll get x^2 -x +2 =0

OpenStudy (anonymous):

right ok I got that.

OpenStudy (callisto):

match the coefficient with a, b,c in the general form

OpenStudy (callisto):

a=1, b=-1, c=2

OpenStudy (callisto):

Then put these value into the quadratic formula, and you'll get x

OpenStudy (callisto):

*values

OpenStudy (anonymous):

so do I really put in the numbers 1 when putting the values into the formula?

OpenStudy (callisto):

put a=1, b=-1, c=2 into the formula

OpenStudy (anonymous):

ok got that...here is what it looks like;